OpenStudy (anonymous):

I'm working on integrating an exact equation and I need help on a certain issue. Is there anyone who has some knowhow on differential equations?

OpenStudy (anonymous):

post ur question plz :)

mathslover (mathslover):

Suraj, please let us kow the quest. So that we can help you accordingly.

OpenStudy (anonymous):

So, the equation I'm dealing with is \[(e^xsin(y) - 2ysin(x))dx + (e^xcos(y) + 2\cos(x))dy = 0\]. I haven't ever had an equation with a dx in there, does that change anything? When I'm finding \[\psi\]?

OpenStudy (anonymous):

thats the type of an exact differential equation, right ? :)

OpenStudy (anonymous):

\[M dx+Ndy=0\]have u seen it before?

OpenStudy (anonymous):

yup just wondering if the dx changes the process at all i have seen that variety, but none of the other homework problems had a dx so

OpenStudy (anonymous):

maybe other hw's are like this\[M+N y'=0\]its the same thing\[M+N \frac{dy}{dx}=0\]multiply by \(dx\)\[Mdx+Ndy =0\]

OpenStudy (anonymous):

k, i hope it helps :)

OpenStudy (anonymous):

OH thank you~! :3

OpenStudy (anonymous):

yw :)